# Efficient Decoding Methods for Language Models on Encrypted Data

### v1

@@ Abstract
Large language models (LLMs) power modern AI applications, but processing sensitive data on untrusted servers raises privacy concerns. Homomorphic encryption (HE) enables computation on encrypted data for secure inference. However, neural text generation requires decoding methods like $\operatorname{argmax}$ and $\operatorname{sampling}$, which are non-polynomial and thus computationally expensive under encryption, creating a significant performance bottleneck. We introduce $\operatorname{CutMax}$, an HE-friendly $\operatorname{argmax}$ algorithm that reduces ciphertext operations compared to prior methods, enabling practical greedy decoding under encryption. We also propose the first HE-compatible nucleus (top-$p$) sampling method, leveraging $\operatorname{CutMax}$ for efficient stochastic decoding with provable privacy guarantees. Both techniques are polynomial, supporting efficient inference in privacy-preserving settings. Moreover, their differentiability facilitates gradient-based sequence-level optimization as a polynomial alternative to straight-through estimators. We further provide strong theoretical guarantees for $\operatorname{CutMax}$, proving it converges globally to a unique two-level fixed point, independent of the input values beyond the identity of the maximizer, which explains its rapid convergence in just a few iterations. Evaluations on realistic LLM outputs show latency reductions of $24\times$$35\times$ over baselines, advancing secure text generation.
Figure 1: Scope: Secure LLM generation over the clients’ encrypted data using HE (model is not necessarily encrypted). Here, standard decoding methods like $\operatorname{argmax}$ and sampling (red) for selecting the next token are not HE-friendly or considered inefficient. We introduce efficient and scalable HE-friendly ways to evaluate them under encryption.
@@ 1 Introduction
Recent advances in LLMs have enabled the development of powerful AI systems capable of generating fluent text at scale (Brown et al., 2020). However, deploying these models in real-world applications often involves sending sensitive user data, such as personal messages or medical records, to remote servers, raising significant privacy concerns (Yao et al., 2024; Yan et al., 2024). Homomorphic encryption (HE) offers a promising solution by allowing computations on encrypted data, ensuring that servers can process queries without accessing their plaintext content (Gentry, 2009). Under HE, users encrypt their inputs, the server performs computations (e.g., running an LLM), and returns an encrypted result that only the user can decrypt. In this paper, we present differentiable and polynomial $\operatorname{argmax}$ and $\operatorname{nucleus}\thinspace\operatorname{sampling}$ algorithms tailored for encrypted LLM decoding. Unlike prior approaches, such as that of Bengio et al. (2013), which use $\operatorname{argmax}$ in the forward pass and estimate gradients in the backward pass via straight-through estimators (STE), our methods are fully differentiable in their plaintext form (without approximations) and remain so under polynomial HE approximations. This differentiability, together with their polynomial form, makes them well suited to AI-privacy settings such as HE; the plaintext formulations are exactly differentiable, and the HE instantiations preserve differentiability via polynomial approximations (see Sect.˜3.2). Most HE schemes, including the CKKS scheme used in this work (Cheon et al., 2017), support only polynomial operations like addition and multiplication. This limitation necessitates the design of HE-friendly algorithms that rely exclusively on such operations, posing a challenge for text generation tasks that involve complex decoding steps. While secure inference on LLMs using HE has been demonstrated (Zimerman et al., 2024a; de Castro et al., 2024; Zhang et al., 2024a), extending these techniques to multi-token text generation remains challenging. Decoding methods like $\operatorname{argmax}$ (for greedy decoding) and $\operatorname{sampling}$ (e.g., nucleus sampling (Holtzman et al., 2019a)) are computationally inefficient or impractical in the encrypted domain due to their reliance on non-polynomial operations.
State-of-the-art homomorphic $\operatorname{argmax}$ implementations follow two comparison-heavy patterns: a tournament tree and an all-to-all league schedule, both realized via deep polynomial approximations of sign (Chakraborty and Zuber, 2022; Crawford et al., 2018; Folkerts and Tsoutsos, 2024; Jovanovic et al., 2022). A detailed illustration of these designs is provided in Figure˜5 (Appendix˜C), which highlights their sequential comparison stages and reliance on costly sign evaluations.
@@
Our contributions address key bottlenecks in secure text generation , as illustrated in Figure˜1, and include: (i) CutMax. An efficient, homomorphic encryption (HE)-friendly $\operatorname{argmax}$ algorithm that enables practical greedy decoding under encryption with fewer ciphertext operations than previous methods. (ii) Encrypted nucleus sampling. The first HE-compatible $\operatorname{nucleus}$ (top-$p$) $\operatorname{sampling}$ method, enables stochastic decoding with provable privacy guarantees, via $\operatorname{CutMax}$. (iii) Theoretical convergence guarantees. A formal proof establishing that $\operatorname{CutMax}$ converges globally to a unique two-level fixed point, providing a rigorous foundation for its rapid convergence in a small, constant number of iterations. (iv) Differentiable decoding primitives. Implementations of $\operatorname{argmax}$ and $\operatorname{nucleus}\thinspace\operatorname{sampling}$ that are real-analytic and fully differentiable in plaintext, leveraging smooth operations like $1/\sqrt{x}$ (and polynomial-only in their HE approximations) to enable exact gradient-based sequence-level training as a theoretically grounded alternative to straight-through estimators. By enabling efficient and accurate multi-token generation under full encryption, our work advances the deployment of privacy-preserving LLMs in real-world settings, bridging a critical gap in secure AI systems.
@@ 2 Background: Homomorphic Argmax
State-of-the-art homomorphic $\operatorname{argmax}$ implementations for LLM decoding rely on comparison-heavy designs, primarily the tournament tree and league schedule, both implemented via deep polynomial approximations of the sign function (Chakraborty and Zuber, 2022; Crawford et al., 2018; Folkerts and Tsoutsos, 2024; Jovanovic et al., 2022; Zhang et al., 2024b, 2025a). In the tournament design, approximately n1n{-}1 pairwise comparisons are organized into $\log_{2}n$ sequential stages, halving the number of candidates per stage until the maximum remains. The league design performs $\binom{n}{2}=\Theta(n^{2})$ comparisons across $n$ sequential rounds, accumulating scores such that the maximum achieves a score of n1n{-}1. Both methods are illustrated in Figure˜5 (Appendix˜C). Despite leveraging CKKS SIMD parallelism to execute multiple comparisons within a round, these approaches incur high multiplicative depth due to repeated sign evaluations, leading to costly bootstrapping and limiting scalability for large vocabularies ($n\sim 10^{5}$). Some works, such as Chakraborty and Zuber (2022); Folkerts and Tsoutsos (2024); Grivet Sébert et al. (2021), use TFHE (Chillotti et al., 2016), restricting input sizes (e.g., $n\leq 256$) due to efficiency constraints. Hybrid approaches combine tournament and league methods to exploit SIMD packing (Iliashenko and Zucca, 2021; Mazzone et al., 2025), but still rely on slow sign approximations. As shown in Table˜3 (Appendix˜B), sign evaluations are significantly deeper and slower than $\operatorname{inverse}$ or $\operatorname{inverse-square-root}$ operations, which our $\operatorname{CutMax}$ algorithm uses instead (Sect.˜3.2). These limitations make prior methods impractical for efficient, privacy-preserving LLM decoding, motivating our polynomial-based approach that eliminates comparisons entirely.
@@ 2.1 HE Preliminaries
HE enables computation on encrypted data without requiring decryption Gentry (2009). Informally, it is defined as follows: Let $\mathcal{R}_{1}(+,\cdot),\mathcal{R}_{2}(\oplus,\odot)$ be two rings for the plaintext and ciphertext spaces, respectively. Given plaintext inputs $m_{1},m_{2}\in\mathcal{R}_{1}$, their encryption $m_{i}]\!]:=\text{Enc}(m_{i})\in\mathcal{R}_{2}$, satisfies the following correctness and homomorphic properties:
@@
where $\epsilon$ is a small noise introduced by the HE scheme, similar to the noise accumulated in floating-point computations. HE is semantically secure and thus every call to $Enc(m_{1})$ results with a new pseudo random ciphertext. Details about the standard threat-model of HE-based applications are provided in Appendix˜A. As HE only supports polynomial operations (additions and multiplications), standard transformer components like $\operatorname{softmax}$ and $\operatorname{layer\ normalization}$ are not directly supported and must be approximated or replaced. Prior-art such as Gilad-Bachrach et al. (2016); Aharoni et al. (2023); Baruch et al. (2022, 2024); Zimerman et al. (2024b, a); de Castro et al. (2024); Zhang et al. (2024a) have proposed polynomial-friendly adaptations for encrypted inference. We focus on the generation process where new challenges arise: decoding operations such as $\operatorname{argmax}$ and $\operatorname{sampling}$ are not polynomial.
@@ 3.1 The CutMax Algorithm
6:  $s^{2}=\frac{1}{n}\sum_{i=1}^{n}(Y^{(t)}_{i}-\mu)^{2}$ $\triangleright$ Variance of $Y^{(t)}$
7:   $Y^{(t+1)}=\left(\frac{Y^{(t)}_{i}-\mu}{c\cdot\sqrt{s^{2}}}+1\right)_{i\in[n]}$ $\triangleright$ Standardization
@@
Figure 2: Illustration of CutMax First Iteration on a Normal Distribution $\operatorname{CutMax}$ algorithm first iteration illustration, applied to the original random variable $X\sim{}\mathrm{Normal}(170,10)$ (a). We first standardize and shift $X$ using $Y=\frac{X-\mu_{X}}{c\sigma_{X}}+1$ with $c=5$ (b), then raise $Y$ to the power $p=3$ (c) or $p=7$ (d) to induce right skew. The resulting distribution $Z=Y^{7}$ (d) has a maximum normalized height $\max\bigl{(}Z/\sum_{i}Z_{i}\bigr{)}=0.0033$, indicating a strong right-tail skew. A vertical black line indicates the mean of every distribution and gray lines indicate the first standard deviation.
@@
The normalization process starts by computing the mean $\mu$ of the current vector state $Y^{(t)}$, the variance $s^{2}$, and the standard deviation $\sigma=\sqrt{s^{2}}$. Subsequently, it scales $Y^{(t)}$ by subtracting the mean and dividing by $c\cdot\sigma$. The choice of $c$ controls how much we “flatten” the distribution before raising the power. A larger $c$ means we divide by a larger number, making the normalized values smaller in magnitude.
@@
The rapid convergence of $\operatorname{CutMax}$ is guaranteed by its behavior as a dynamical system. As formally proven in Appendix E, one $\operatorname{CutMax}$ inner update followed by normalization strictly reduces the dispersion among all non-maximum entries in standardized space, driving them toward their common mean. In particular, the non-top residual dispersion contracts geometrically by an explicit factor $\kappa$ (see (14)) whenever $c>\sqrt{n-1}$. Consequently, any initial vector with a unique maximum evolves toward a two–level state in which all non-maximum elements are equal, and the normalized iterates converge to a unique two-level fixed point that depends only on (n,c,p)(n,c,p)—independent of the input beyond the identity of its maximal coordinate (cf. Theorem˜E.1).
Beyond this theoretical contraction, $\operatorname{CutMax}$ also prunes much more aggressively than comparison-based schemes. Unlike the tournament $\operatorname{argmax}$, which discards only half of the ${Y}^{(t)}$ values per stage, $\operatorname{CutMax}$ zeroes out a large majority each round: empirically, after iteration $t$, the empirical CDF at the mean satisfies $\mathrm{CDF}_{Y^{(t)}}(\mu^{(t)})\gg\tfrac{1}{2}$, so far more than half of the $Y^{(t)}_{i}$ fall below the mean and are driven toward zero by the subsequent odd-power step (Figure˜3).
As established by Theorem˜E.1, these properties guarantee global convergence to the fixed point under mild conditions; for the practical hyperparameters we use, the limit is nearly one-hot. This explains why only a small, nearly constant number of iterations (e.g., $T\!\leq\!3\text{\textendash}4$) suffices in practice, regardless of the initial logit distribution (Sect.˜4).
@@ 3.2 HE-Friendly CutMax
Building on the HE preliminaries in Sect.˜2, we adapt $\operatorname{CutMax}$ to be fully polynomial by approximating $1/\sigma$ (Line 5, Algorithm˜1) and $1/\sum_{j}Y^{(T)}_{j}$ (Line 7) using Goldschmidt methods Goldschmidt (1964) (details in Sect.˜B.1). This yields Algorithm˜2, the FHE variant. Notably, having a polynomial algorithm is not enough, we also need to ensure that no overflows or precision issues compromise the correctness of the results. Specifically, we must prove the following: (i) No overflows. Let $B_{\text{CKKS}}$ be the bound configured by the scheme on the unencrypted inputs, beyond which an overflow may occur. Then, all intermediate values produced by the algorithm must remain below $B_{\text{CKKS}}$. (ii) No approximation errors. The $\operatorname{InvSqr}$ and $\operatorname{Inv}$ approximations assume inputs lie within a certain range, e.g., $\frac{1}{2^{n}},1]$ for some small $n$. If an input $x$ falls in a wider range $[a,b]$, we instead compute $\frac{\operatorname{InvSqr}(\frac{x}{b})}{b}$. We must show that $\frac{a}{b}>\frac{1}{2^{n}}$; otherwise, the approximation may yield incorrect results. Sect.˜2 reviewed the tournament method that requires $O(n)$ comparisons in $\log n$ sequential stages and the league method that requires $n^{2}$ comparisons in $n$ sequential stages. In contrast, $\operatorname{CutMax}$ requires only $T$ $\operatorname{InvSqr}$ and 1 $\operatorname{Inv}$ operations. Table˜3 in Appendix˜B (taken from Moon et al. (2024a)) shows that the $\operatorname{InvSqr}$ implementation is $1.5\text{--}2.1\times$ faster than the sign implementation used in all prior art, at the same error level. In addition, $T$ is weakly-dependent on $n$ and much smaller. Consequently, $\operatorname{CutMax}$ outperforms prior-art. Unfortunately, the situation is not that simple. Modern HE schemes support single instruction multiple data (SIMD) operations, meaning a ciphertext can encrypt a vector of $s$ elements instead of just a single element. In this case, multiplications ($\odot$), additions ($\oplus$), and subtractions ($\ominus$) are applied element-wise homomorphically over the vector. Furthermore, a homomorphic rotation by $\ell<s$ positions is also available (Cheon et al., 2017).
@@
6:  $s^{2}]\!]=\frac{1}{n}\operatorname{RotAndSum}\left(([\![Y^{(t)}]\!]\ominus[\![\mu]\!])^{2}\right)$
7:  $\sigma^{-1}]\!]=\operatorname{InvSqr}([\![s^{2}]\!])$
@@
We defer the discussion of how SIMD impacts latency to Appendix˜D, and here we complete the picture with an implementation of $\operatorname{CutMax}$ over HE. Algorithm˜2 is the HE approximation variant of $\operatorname{CutMax}$, where the input vector $X$ is encrypted. For brevity, we assume $n=a\cdot s$, $a\in\mathbb{Z}_{+}$, and write the algorithm so that modern compilers such as HELayers Aharoni et al. (2023) can automatically implement it. Lines 3, 4, and 8 of Algorithm˜1 are efficiently computed in Algorithm˜2 using the $\operatorname{RotAndSum}$ method (see Adir et al. (2024)), which takes an encrypted vector $v=(v_{1},\ldots,v_{s})$ and returns a vector of $s$ elements, each containing the sum $\sum v_{i}$. This requires $\log s$ rotations and additions. Multiplying by a plaintext scalar ($\frac{1}{n}$ at lines 3,4 and $\frac{1}{c}$ at line 6) is considered a cheap HE operation as well as the homomorphic subtractions at lines 4 and 6. Raising to the power of 2 (Line 4) and $p$ (Line 6) require 1 and $\log p$ multiplications, respectively. Finally, the heaviest operations used are the $\operatorname{InvSqr}$ (Line 5) and $\operatorname{Inv}$ (Line 8) that use the Goldschmidt approximations (Sect.˜B.1). Denote by $M_{\operatorname{Inv}}$ (resp. $D_{\operatorname{Inv}}$) and $M_{\operatorname{InvSqr}}$ (resp. $D_{\operatorname{InvSqr}}$) the number of multiplications (resp. multiplication depth) required by the $\operatorname{Inv}$ and $\operatorname{InvSqr}$ approximations over one ciphertext, respectively. Then $\operatorname{CutMaxHE}$ executes

### v2

@@ Abstract
Large language models (LLMs) power modern AI applications, but processing sensitive data on untrusted servers raises privacy concerns. Homomorphic encryption (HE) enables computation on encrypted data for secure inference. However, neural text generation requires decoding methods like $\operatorname{argmax}$ and $\operatorname{sampling}$, which are non-polynomial and thus computationally expensive under encryption, creating a significant performance bottleneck. We introduce $\operatorname{CutMax}$, an HE-friendly $\operatorname{argmax}$ algorithm that reduces ciphertext operations compared to prior methods, enabling practical greedy decoding under encryption. We also propose the first HE-compatible nucleus (top-$p$) sampling method, leveraging $\operatorname{CutMax}$ for efficient stochastic decoding with provable privacy guarantees. Both techniques are polynomial, supporting efficient inference in privacy-preserving settings. Moreover, their differentiability facilitates gradient-based sequence-level optimization as a polynomial alternative to straight-through estimators. We further provide strong theoretical guarantees for $\operatorname{CutMax}$, proving its convergence via exponential amplification of the gap ratio between the maximum and runner-up elements. Evaluations on realistic LLM outputs show latency reductions of $24\times$$35\times$ over baselines, advancing secure text generation.
Figure 1: Scope: Secure LLM generation over the clients’ encrypted data using HE (model is not necessarily encrypted). Here, standard decoding methods like $\operatorname{argmax}$ and $\operatorname{sampling}$ (red) for selecting the next token are not HE-friendly or considered inefficient. We introduce efficient and scalable HE-friendly methods to evaluate them under encryption.
@@ 1 Introduction
Recent advances in LLMs have enabled the development of powerful AI systems capable of generating fluent text at scale (brown2020language). However, deploying these models in real-world applications often involves sending sensitive user data, such as personal messages or medical records, to remote servers, raising significant privacy concerns (yao2024survey; yan2024protecting). Despite major progress in efficient encrypted inference, extending these methods to generative LLMs remains an open challenge due to the non-polynomial nature of decoding operations. Homomorphic encryption (HE) offers a promising solution by allowing computations on encrypted data, ensuring that servers can process queries without accessing their plaintext content (Gentry2009). Under HE, users encrypt their inputs, the server performs computations (e.g., running an LLM), and returns an encrypted result that only the user can decrypt.
In this paper, we present differentiable and polynomial $\operatorname{argmax}$ and $\operatorname{nucleus}\thinspace\operatorname{sampling}$ algorithms tailored for encrypted LLM decoding. Unlike prior approaches, such as that of Bengio2013EstimatingOP, which use $\operatorname{argmax}$ in the forward pass and estimate gradients in the backward pass via straight-through estimators (STE), our methods are fully differentiable in their plaintext form (without approximations) and remain so under polynomial HE approximations. This differentiability, together with their polynomial form, makes them well suited to AI-privacy settings such as HE; the plaintext formulations are exactly differentiable, and the HE instantiations preserve differentiability via polynomial approximations (see Sect.˜3.2). Most HE schemes, including the CKKS scheme used in this work (ckks2017), support only polynomial operations like addition and multiplication. This limitation necessitates the design of HE-friendly algorithms that rely exclusively on such operations, posing a challenge for text generation tasks that involve complex decoding steps. While secure inference on LLMs using HE has been demonstrated (zimerman2024power; Castro_2024; transformerArgmax), extending these techniques to multi-token text generation remains challenging. Decoding methods like $\operatorname{argmax}$ (for greedy decoding) and $\operatorname{sampling}$ (e.g., $\operatorname{nucleus}\thinspace\operatorname{sampling}$ (holtzman2019curious)) are computationally inefficient or impractical in the encrypted domain due to their reliance on non-polynomial operations.
State-of-the-art homomorphic $\operatorname{argmax}$ implementations follow two comparison-heavy patterns: a tournament tree and an all-to-all league schedule, both realized via deep polynomial approximations of sign (HeCompArgmaxTournament; LRArgmaxTournament; Tyche; pheonix2022). A detailed illustration of these designs is provided in Figure˜6 (Appendix˜C), which highlights their sequential comparison stages and reliance on costly sign evaluations.
@@
Our contributions address key bottlenecks in secure text generation , as illustrated in Figure˜1, and include: (i) CutMax. An efficient, homomorphic encryption (HE)-friendly $\operatorname{argmax}$ algorithm that enables practical greedy decoding under encryption with fewer ciphertext operations than previous methods. (ii) Encrypted nucleus sampling. The first HE-compatible $\operatorname{nucleus}$ (top-$p$) $\operatorname{sampling}$ method, enables stochastic decoding with provable privacy guarantees, via $\operatorname{CutMax}$. (iii) Theoretical convergence guarantees. A formal proof establishing that $\operatorname{CutMax}$ converges via exponential amplification of the gap ratio between the maximum and runner-up elements, providing a rigorous foundation for its rapid convergence in a small number of iterations, weakly dependent on the array size. (iv) Differentiable decoding primitives. Implementations of $\operatorname{argmax}$ and $\operatorname{nucleus}\thinspace\operatorname{sampling}$ that are real-analytic and fully differentiable in plaintext, leveraging smooth operations like $1/\sqrt{x}$ (and polynomial-only in their HE approximations) to enable exact gradient-based sequence-level training as a theoretically grounded alternative to straight-through estimators. By enabling efficient and accurate multi-token generation under full encryption, our work advances the deployment of privacy-preserving LLMs in real-world settings, bridging a critical gap in secure AI systems.
@@ 2 Background: Homomorphic Argmax
State-of-the-art homomorphic $\operatorname{argmax}$ implementations for LLM decoding rely on comparison-heavy designs, primarily the tournament tree and league schedule, both implemented via deep polynomial approximations of the sign function (HeCompArgmaxTournament; LRArgmaxTournament; Tyche; pheonix2022; Zhang_2024; zhang2025secpesecurepromptensembling). In the tournament design, approximately n1n{-}1 pairwise comparisons are organized into $\log_{2}n$ sequential stages, halving the number of candidates per stage until the maximum remains. The league design performs $\binom{n}{2}=\Theta(n^{2})$ comparisons across $n$ sequential rounds, accumulating scores such that the maximum achieves a score of n1n{-}1. Both methods are illustrated in Figure˜6 (Appendix˜C). Despite leveraging CKKS SIMD parallelism to execute multiple comparisons within a round, these approaches incur high multiplicative depth due to repeated sign evaluations, leading to costly bootstrapping and limiting scalability for large vocabularies ($n\sim 10^{5}$). Some works, such as HeCompArgmaxTournament; Tyche; grivet2021speed, use TFHE (tfhe), restricting input sizes (e.g., $n\leq 256$) due to efficiency constraints. Hybrid approaches combine tournament and league methods to exploit SIMD packing (BGVCompArgmaxCombined; tournamentandleague), but still rely on slow sign approximations. As shown in Table˜3 (Appendix˜B), sign evaluations are significantly deeper and slower than $\operatorname{inverse}$ or $\operatorname{inverse-square-root}$ operations, which our $\operatorname{CutMax}$ algorithm uses instead (Sect.˜3.2). These limitations make prior methods impractical for efficient, privacy-preserving LLM decoding, motivating our polynomial-based approach that eliminates comparisons entirely.
@@ 2.1 HE Preliminaries
HE enables computation on encrypted data without requiring decryption Gentry2009. Informally, it is defined as follows: Let $\mathcal{R}_{1}(+,\cdot),\mathcal{R}_{2}(\oplus,\odot)$ be two rings for the plaintext and ciphertext spaces, respectively. Given plaintext inputs $m_{1},m_{2}\in\mathcal{R}_{1}$, their encryption $m_{i}]\!]:=\text{Enc}(m_{i})\in\mathcal{R}_{2}$, satisfies the following correctness and homomorphic properties:
@@
where $\epsilon$ is a small noise introduced by the HE scheme, similar to the noise accumulated in floating-point computations. HE is semantically secure and thus every call to $Enc(m_{1})$ results with a new pseudo random ciphertext. Details about the standard threat-model of HE-based applications are provided in Appendix˜A. As HE only supports polynomial operations (additions and multiplications), standard transformer components like $\operatorname{Softmax}$ and $\operatorname{LayerNormalization}$ are not directly supported and must be approximated or replaced. Prior-art such as Dowlin_2016; helayers; baruch2022methodology; baruch2024polynomial; zimerman2024converting; zimerman2024power; Castro_2024; transformerArgmax have proposed polynomial-friendly adaptations for encrypted inference. We focus on the generation process where new challenges arise: decoding operations such as $\operatorname{argmax}$ and $\operatorname{sampling}$ are not polynomial.
@@ 3.1 The CutMax Algorithm
6:  $\sigma^{2}=\frac{1}{n}\sum_{i=1}^{n}(Y^{(t)}_{i}-\mu)^{2}$ $\triangleright$ Variance of $Y^{(t)}$
7:   $Y^{(t+1)}=\left(\frac{Y^{(t)}_{i}-\mu}{c\cdot\sqrt{\sigma^{2}}}+1\right)_{i\in[n]}$ $\triangleright$ Standardization
@@
Figure 2: Illustration of CutMax First Iteration on a Normal Distribution $\operatorname{CutMax}$ algorithm first iteration illustration, applied to the original random variable $X\sim{}\mathrm{Normal}(170,10)$ (a). We first standardize and shift $X$ using $Y=\frac{X-\mu_{X}}{c\sigma_{X}}+1$ with $c=5$ (b), then raise $Y$ to the power $p=3$ (c) or $p=7$ (d) to induce right skew. The resulting distribution $Z=Y^{7}$ (d) has a maximum normalized height $\max\bigl(Z/\sum_{i}Z_{i}\bigr)=0.0033$, indicating a strong right‐tail skew. A vertical black line indicates the mean of every distribution and gray lines indicate the first standard deviation.
@@
The normalization process starts by computing the mean $\mu$ of the current vector state $Y^{(t)}$, the variance $\sigma^{2}$, then the inverse standard deviation $\frac{1}{\sqrt{\sigma^{2}}}$. Subsequently, it scales $Y^{(t)}$ by subtracting the mean and multiplying by $\frac{1}{c\cdot\sigma}$. The choice of $c$ controls how much we “flatten” the distribution before raising the power. A larger $c$ means we divide by a larger number, making the normalized values smaller in magnitude.
@@
As formally proven in Appendix E, each $\operatorname{CutMax}$ iteration unconditionally grows the gap between the maximum and runner-up elements, yielding exponential amplification of their ratio and ensuring convergence under a mild $\delta$-gap condition on the input. Beyond this theoretical contraction, $\operatorname{CutMax}$ also prunes much more aggressively than comparison-based schemes. Unlike the tournament $\operatorname{argmax}$, which discards only half of the ${Y}^{(t)}$ values per stage, $\operatorname{CutMax}$ zeroes out a large majority each round: empirically, after iteration $t$, the empirical CDF at the mean satisfies $\mathrm{CDF}_{Y^{(t)}}(\mu^{(t)})\gg\tfrac{1}{2}$, so far more than half of the $Y^{(t)}_{i}$ fall below the mean and are driven toward zero by the subsequent odd-power step (Figure˜3).
As established by Theorem˜E.10, these properties guarantee global convergence to the fixed point under mild conditions; for the practical hyperparameters we use, the limit is nearly one-hot. This explains why only a small, nearly constant number of iterations (e.g., $T\!\leq\!3\text{–}4$) suffices in practice, regardless of the initial logit distribution (Sect.˜4).
@@ 3.2 HE-Friendly CutMax
Building on the HE preliminaries in Sect.˜2, we adapt $\operatorname{CutMax}$ to be fully polynomial by approximating $1/\sigma$ (Line 5, Algorithm˜1) and $1/\sum_{j}Y^{(T)}_{j}$ (Line 7) using Goldschmidt methods goldschmidt (details in Sect.˜B.1). This yields Algorithm˜2, the FHE variant. Notably, having a polynomial algorithm is not enough, we also need to ensure that no overflows or precision issues compromise the correctness of the results. Specifically, we must prove the following:
(i) No overflows. Let $B_{\text{CKKS}}$ be the bound configured by the scheme on the unencrypted inputs, beyond which an overflow may occur. Then, all intermediate values produced by the algorithm must remain below $B_{\text{CKKS}}$.
(ii) No approximation errors. The $\operatorname{InvSqr}$ and $\operatorname{Inv}$ approximations assume inputs lie within a certain range, e.g., $\frac{1}{2^{n}},1]$ for some small $n$. If an input $x$ falls in a wider range $[a,b]$, we instead compute $\frac{\operatorname{InvSqr}(\frac{x}{b})}{b}$. We must show that $\frac{a}{b}>\frac{1}{2^{n}}$; otherwise, the approximation may yield incorrect results. Sect.˜2 reviewed the tournament method that requires $O(n)$ comparisons in $\log n$ sequential stages and the league method that requires $n^{2}$ comparisons in $n$ sequential stages. In contrast, $\operatorname{CutMax}$ requires only $T$ $\operatorname{InvSqr}$ and 1 $\operatorname{Inv}$ operations. Table˜3 in Appendix˜B (taken from asor) shows that the $\operatorname{InvSqr}$ implementation is $1.5\text{--}2.1\times$ faster than the sign implementation used in all prior art, at the same error level. In addition, $T$ is weakly-dependent on $n$ and much smaller. Consequently, $\operatorname{CutMax}$ outperforms prior-art. Unfortunately, the situation is not that simple. Modern HE schemes support single instruction multiple data (SIMD) operations, meaning a ciphertext can encrypt a vector of $s$ elements instead of just a single element. In this case, multiplications ($\odot$), additions ($\oplus$), and subtractions ($\ominus$) are applied element-wise homomorphically over the vector. Furthermore, a homomorphic rotation by $\ell<s$ positions is also available (ckks2017).
@@
6:  $\sigma^{2}]\!]=\frac{1}{n}\operatorname{RotAndSum}\left(([\![Y^{(t)}]\!]\ominus[\![\mu]\!])^{2}\right)$
7:  $\sigma^{-1}]\!]=\operatorname{InvSqr}([\![\sigma^{2}]\!])$
@@
We defer the discussion of how SIMD impacts latency to Appendix˜D, and here we complete the picture with an implementation of $\operatorname{CutMax}$ over HE. Algorithm˜2 is the HE approximation variant of $\operatorname{CutMax}$, where the input vector $X$ is encrypted. For brevity, we assume $n=a\cdot s$, $a\in\mathbb{Z}_{+}$, and write the algorithm so that modern compilers such as HELayers helayers can automatically implement it. Lines 3, 4, and 8 of Algorithm˜1 are efficiently computed in Algorithm˜2 using the $\operatorname{RotAndSum}$ method (see he4ds), which takes an encrypted vector $v=(v_{1},\ldots,v_{s})$ and returns a vector of $s$ elements, each containing the sum $\sum v_{i}$. This requires $\log s$ rotations and additions. Multiplying by a plaintext scalar ($\frac{1}{n}$ at lines 3,4 and $\frac{1}{c}$ at line 6) is considered a cheap HE operation as well as the homomorphic subtractions at lines 4 and 6. Raising to the power of 2 (Line 4) and $p$ (Line 6) require 1 and $\log p$ multiplications, respectively. Finally, the heaviest operations used are the $\operatorname{InvSqr}$ (Line 5) and $\operatorname{Inv}$ (Line 8) that use the Goldschmidt approximations (Sect.˜B.1). Denote by $M_{\operatorname{Inv}}$ (resp. $D_{\operatorname{Inv}}$) and $M_{\operatorname{InvSqr}}$ (resp. $D_{\operatorname{InvSqr}}$) the number of multiplications (resp. multiplication depth) required by the $\operatorname{Inv}$ and $\operatorname{InvSqr}$ approximations over one ciphertext, respectively. Then $\operatorname{CutMaxHE}$ executes